OpenStudy (anonymous):

Which recursive formula can be used to generate the sequence shown, where f(1) = 9.6 and n > 1? 9.6, –4.8, 2.4, –1.2, 0.6, ... f(n + 1) = –0.5f(n) f(n + 1) = 0.5f(n) f(n + 1) = f(0.5n) f(n + 1) = f(–0.5n)

OpenStudy (welshfella):

plug in 1 for n and 9.6 for f(n) and see which one of the options fit

OpenStudy (welshfella):

so for example, try th first one:- f(1 +1) =f(2) = -0.48 should equal -(0.5)* 9.6

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

what does -0.5 * 9.6 work out to ?

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

f(n+1)=-0.5f(n)

OpenStudy (anonymous):

f(n+1)=-0.5f(n)

OpenStudy (anonymous):

thank you